OpenStudy (anonymous):

6+9 square root of 5/3

OpenStudy (anonymous):



OpenStudy (anonymous):

do you do 6+9 divided by 3 which is 5 and then multiply square root 5?

OpenStudy (anonymous):

Simply following PEMDAS will get you the answer here.

OpenStudy (anonymous):

as you have it written you cannot add 6 to the 9 because the square root is attached to the 9 but you can divide 6 by 3 and 9 by 3 so you end up with 2+3sqrt(5)

OpenStudy (zzr0ck3r):

\(\frac{a+b}{c}=\frac{a}{c}+\frac{b}{c}\)

OpenStudy (zzr0ck3r):

\(\frac{6+9\sqrt{5}}{3}=\frac{6}{3}+\frac{9\sqrt{5}}{3}=\frac{6}{3}+\frac{9}{3}\sqrt{5}\)

OpenStudy (zzr0ck3r):

\(\huge = ?\)

OpenStudy (anonymous):

2+3 square root 5

OpenStudy (anonymous):

because you cant simplify 5 further right?

OpenStudy (zzr0ck3r):

correct, never if the number is prime.

OpenStudy (anonymous):

thanks so much!

OpenStudy (zzr0ck3r):

and many times not, even if it is not prime, like sqrt(15)

OpenStudy (zzr0ck3r):

np